# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

### v38

@@ 4.2 Sufficient Statistics
(i) The noise-adjusted projection coefficient of the market maker’s order flow $\omega$ onto $\widetilde{W}(\,\cdot\,|s_{k})$ is
$\Pi_{mm}(\omega,s_{k};\widetilde{W})=\sum_{i}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}}.$ (14)
(ii) The noise-adjusted projection coefficient of an insider portfolio $W(\,\cdot\,)$ onto $\widetilde{W}(\,\cdot\,|s_{k})$ is
@@ 5.1 First-Order Condition
In trading to exploit his private information, the insider balances each security’s marginal payoff against its marginal cost. This cost comprises, first, the security’s current price (conditional on no price impact) and, second, its price impact across securities. Theorem 5.1 makes this intuition precise in the general setting.
@@
The insider’s marginal utility from adding a marginal portfolio $v$ to a portfolio $W$ decomposes into
$\underbrace{\partial\!\kern 0.80002ptJ(v;W)}_{\text{marginal utility}}=\underbrace{\partial\!\kern 0.80002ptJ_{p}(v;W)}_{\text{marginal payoff}}-\underbrace{\big(\,\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!AD}(v;W)+\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)\,\big)}_{\text{marginal cost}}$
where the terms are defined as follows:
@@
Therefore, any optimal portfolio $W$ must satisfy the first-order condition
$\partial\!\kern 0.80002ptJ_{p}(v;W)=\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!AD}(v;W)+\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)\quad\mbox{for all marginal portfolio $v$}.$ (20)
Conditional on signal $s_{k}$, the marginal payoff of a security $x_{i}$ is its true payoff $\eta(x_{i}|s_{k})$ known to the insider. Indeed, Equation (17) shows that the marginal payoff functional $\partial\!\kern 0.80002ptJ_{p}(\,\cdot\,;W)$ is identified with the true payoff distribution $\eta(\,\cdot\,|s_{k})$. The first-order condition (20) equalizes this marginal payoff $\partial\!\kern 0.80002ptJ_{p}(v;W)$ with the marginal cost consisting of two terms:
[…]
  • Price Impact Term $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)$: This term is the cost incurred due to the market maker adjusting prices across markets in response to the insider adding marginal portfolio $v$.
5.2 The Price Impact Term
From the expression of $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)$ in (19), it follows that the price impact of the portfolio $v$ on security $x_{i}$, conditional on $\omega$, is
${\operatorname{\mathbb{Cov}}}\left(\eta(x_{i}|\,\cdot\,),\Pi_{insider}(v,\,\cdot\,;\widetilde{W})\,\bigg|\,\omega\right).$ (21)
@@ 5.2 The Price Impact Term
Given market maker belief $\widetilde{W}$, the cross price impact of $x_{j}$ on $x_{i}$, conditional on $\omega$, is
@@
The expected cross price impact is given by
@@
The intuition underlying the cross price impact characterization in Equation (22) is straightforward: demand for $x_{j}$ increases the price of $x_{i}$ if the market maker believes that a high demand for $x_{j}$ correlates with a high payoff of $x_{i}$—in other words, if he believes there is positive correlation between the insider’s demand $\widetilde{W}(x_{j}|\,\cdot\,)$ for $x_{j}$ and the payoff $\eta(x_{i}|\,\cdot\,)$ of $x_{i}$ across signals, and vice versa. A greater noise trading intensity $\sigma_{j}$ for $x_{j}$ reduces this price impact.
Example 5.3.
Suppose a security $x_{j}$ pays off only under signal $s_{k_{j}}$ and not any other signal. Suppose further that the market maker reasonably believes that the insider’s demand for $x_{j}$ is positive under $s_{k_{j}}$ but zero otherwise. Consider the following two scenarios for another security $x_{i}$:
  1. (i)
    (Positive Price Impact) Suppose $x_{i}$ also pays off only under $s_{k_{j}}$, the same signal that triggers payoff for $x_{j}$. Then $x_{j}$ has a positive price impact on $x_{i}$: a strong demand for $x_{j}$ suggests to the market maker that $x_{i}$ is also likely to pay off, leading him to raise the price of $x_{i}$. Indeed, here the payoff $\eta(x_{i}|\,\cdot\,)$ and $\widetilde{W}(x_{j}|\,\cdot\,)$ are positively correlated, both being positive at $s_{k_{j}}$ and zero otherwise. Therefore, by Corollary 5.2, $\frac{\partial}{\partial W(x_{j})}P(x_{i},\omega\kern 0.44998pt;\widetilde{W})>0$.
  2. (ii)
    (Negative Price Impact) Suppose $x_{i}$ only pays off under a signal different from $s_{k_{j}}$. By Corollary 5.2, $x_{j}$ has a negative price impact on $x_{i}$. The intuition is the flip side of scenario (i).
@@ 5.3 No-Arbitrage
No-arbitrage is a general necessary condition for the existence of equilibrium. We now formulate this condition for our setting. Here, once again, the projection coefficient profile $\Pi_{insider}$ between the insider’s trades and the market maker’s belief naturally occurs. Given market maker belief $\widetilde{W}$, we define the corresponding zero price impact portfolios as portfolios $W$ for which $\Pi_{insider}(W,\,\cdot\,;\widetilde{W})=0$. Intuitively, such a portfolio causes zero price impact because it is orthogonal to the market maker’s belief about the insider’s trading strategy. These zero price impact portfolios must yield a zero payoff for the insider; otherwise, they would present arbitrage opportunities. This is formally stated in Theorem 5.4.
@@
(No-Arbitrage)
For any given market maker belief, an optimal portfolio for the insider can exist only if the corresponding zero price impact portfolios yield zero payoff for the insider. Otherwise, the insider could obtain unbounded utility.
@@ 6 Equilibrium
(w.l.o.g.)   The market maker has a uniform prior on $S$.
Assumption 1 is made without loss of generality. Any other prior $\pi_{0}$ can be reinterpreted as if it were the equivalent case of the uniform prior $\pi_{u}$ and modified security payoffs $\{\frac{\pi_{0}(s_{k})}{\pi_{u}(s_{k})}\eta(\,\cdot\,|s_{k})\}_{s_{k}\in S}$. [^15] Our proofs go through for any payoff function $\eta(\,\cdot\,|\,\cdot\,)\colon X\times S\rightarrow\mathbb{R}$, including the modified payoffs. Once equilibrium is established under the uniform prior, this equivalence yields the comparative statics for varying the prior. For instance, if two securities both yield higher payoffs under some $s_{k}$ compared to the other signals, then their cross price impact increases with the prior probability of $s_{k}$. This generalizes from the simplified setting in Section 2 (see footnote 9).
@@ 6.1 Canonical Game
In equilibrium, the insider’s zero payoff portfolios must coincide with the zero price impact portfolios. Zero payoff implies zero price impact because the market maker’s equilibrium belief is correct per Definition 1. Conversely, by the no-arbitrage condition of Theorem 5.4, zero price impact implies zero payoff.
The zero payoff portfolios are those orthogonal to the linear span of $\{\eta(\,\cdot\,|s_{k})\}_{k}$—any portfolio $W$ satisfying $\sum\limits_{i}W(x_{i})\eta(x_{i}|s_{k})=0$ for all $s_{k}$ yields zero payoff for the insider. Thus, the insider’s payoff-relevant portfolios lie within the linear span of $\{\eta(\,\cdot\,|s_{k})\}_{k}$. We state this result as Proposition 6.1(i). On the other hand, the market maker’s noise-adjusted equilibrium beliefs must contain this span; otherwise, zero price impact allows arbitrage by the insider. This is Proposition 6.1(ii).
@@
(ii) Conversely, each $\eta(\,\cdot\,|s_{k})$ must lie in the linear span of $\{\frac{W^{*}(\,\cdot\,|s_{l})}{\sigma_{\,\cdot\,}^{2}}\}_{l}$ to prevent arbitrage opportunities.
@@
$\widetilde{W}(\,\cdot\,|s_{k})=\sum_{l=1}^{K}\tilde{d}^{(k)}_{l}\eta(\,\cdot\,|s_{l}),\;\tilde{d}^{(k)}\in\mathbb{R}^{K}\;\mbox{ for }1\leq k\leq K,$
by the matrix $\widetilde{D}=[\tilde{d}^{(1)}\;\cdots\;\tilde{d}^{(K)}]\in\mathbb{R}^{K\times K}$.
@@
Define the information intensity matrix $\mathbf{L}\in\mathbb{R}^{K\times K}$ of the trading game by [^16] In other words, $\bf L$} is the positive-semidefinite square root of $\left[\sum\limits_{i}\frac{\eta(x_{i}|s_{k})\eta(x_{i}|s_{l})}{\sigma_{i}^{2}}\right]_{1\leq k,l\leq K}$. It is positive definite when the $\eta(\,\cdot\,|s_{k})$’s are linearly independent.
${\bf L}^{2}=\left[\sum_{i}\frac{\eta(x_{i}|s_{k})\eta(x_{i}|s_{l})}{\sigma_{i}^{2}}\right]_{1\leq k,l\leq K}\;\;\mbox{ and }\;\;{\bf L}^{T}={\bf L}.$
$\bf L$} generalizes the noise-adjusted variation $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ from the single-asset setting (6). A larger $\bf{L}$ means greater noise-adjusted (co-)variations in payoffs across securities and signals, reflecting higher information intensity. [^17] Here, $\bf L$} is ordered based on positive-semidefiniteness. This is a partial order. Unlike the single-asset setting, the information intensity across securities and signals is not always directly comparable between two specifications.
By Proposition 6.1, the columns of $\bf L$} span the expansion coefficients of the insider’s payoff-relevant portfolios in equilibrium. Therefore, we can apply the canonical transformation
$d\mapsto\hat{d}={\bf L}d.$ (24)
Under this transformation, a market maker belief $\widetilde{D}$ transforms as
$\widehat{D}={\bf L}\widetilde{D}=[\hat{d}^{(1)}\cdots\hat{d}^{(K)}].$ (25)
Canonical Transformation: Intuition
The canonical transformation (24) normalizes the trading game by the insider’s information intensity. To see this intuition clearly, return to the simplified scenario of Section 2 where the payoff distributions $\eta(\,\cdot\,|s_{k})$ do not overlap across states. In this scenario, because there is zero covariation of payoffs across signals, the information intensity matrix $\bf L$} is diagonal. The $k$-th diagonal entry of $\bf L$} is $c_{k}^{2}=\sum\limits_{i}\frac{\eta(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}$, which is the noise-adjusted payoff variation across securities conditional on $s_{k}$. The canonical transformation rescales each payoff distribution $\eta(\,\cdot\,|s_{k})$ by the factor $\frac{1}{c_{k}}$, thus converting $\bf L$} to the identity matrix. In other words, this transformation normalizes the payoff variations across signals to unity. [^18] We note again that our proofs apply to any positive function $\eta(\,\cdot\,|\,\cdot\,)\colon X\times S\rightarrow\mathbb{R}$, including the rescaled payoffs considered here. This normalization yields an isomorphic game that is invariant with respect to $\eta(\,\cdot\,|\,\cdot\,)$ and $(\sigma_{i})$ and symmetric across signals.
@@
Under Assumption 1, the Bayesian trading game between the insider and market maker is isomorphic to the canonical game, a pseudo-trading game defined as follows:
  • There are markets for pseudo-securities $k=1,\cdots,K$.
  • Conditional on signal $s_{k}$, pseudo-security $k$ has payoff of $1$ while all others have payoffs of 0.
  • The insider submits orders $\hat{d}\in\mathbb{R}^{K}$ for the pseudo-securities, and the market maker receives order flow $\hat{\omega}=\hat{d}+\widehat{N}$, where $\widehat{N}_{k}$ are i.i.d. $\mathcal{N}(0,1)$ noise trades for $k=1,\cdots,K$.
  • The market maker has a belief $\widehat{D}$ as specified in (25). Upon receiving $\hat{\omega}$, he updates the posterior probability of each signal $s_{k}$ (which is also the pseudo-security $k$ price) to
    $\hat{\pi}_{1}(k|\hat{\omega}\kern 0.44998pt;\widehat{D})\propto e^{(\hat{d}^{(k)})^{T}\hat{\omega}-\frac{1}{2}(\hat{d}^{(k)})^{T}\hat{d}^{(k)}}.$ (26)
[…]
Equilibrium in Canonical Game
In the canonical game, an equilibrium is specified by a $K\times K$ strategy matrix $D^{*}=[\delta^{(1)}\;\cdots\;\delta^{(K)}]$ such that, under the same market maker belief $D^{*}$, each strategy $\delta^{(k)}\in\mathbb{R}^{K}$ solves the insider’s problem (27) conditional on $s_{k}$, i.e.,
$\delta^{(k)}\in\operatorname*{argmax}_{\hat{d}\,\in\,\mathbb{R}^{K}}J(\hat{d}|k\kern 0.44998pt;D^{*}),\;\mbox{for each}\;k=1,\cdots,K.$ (28)
Corollary 6.3.
Given the isomorphism between the two games, an equilibrium $D^{*}$ of the canonical game maps to the equilibrium of the original trading game (see Definition 1) specified by
[…]
We refer to the matrix ${\bf L}^{-1}D^{*}$ as the canonical form of the original game equilibrium (29).
Remark 6.4.
When the $\eta(\,\cdot\,|s_{k})$’s are linearly dependent, ${\bf L}^{-1}$ is interpreted as the pseudoinverse of $\bf L$}. Compared to the case of linearly independent $\eta(\,\cdot\,|s_{k})$’s, the only caveat is that the informed demand portfolio $W^{*}(\,\cdot\,|s_{k})$ of (29) is unique only up to its payoff, but not the allocations of its constituent $\eta(\,\cdot\,|s_{l})$’s.
Insider FOC in Canonical Game
Let $p\in[0,1]^{I}$ denote the random vector (26) of the market maker’s posterior in the canonical game (with the dependence on $\hat{\omega}$ and $\widehat{D}$ understood) and $\mbox{diag}(p)\in\mathbb{R}^{K\times K}$ denote the matrix with $p$ along the diagonal and zeros elsewhere. Then, conditional on observing $s_{k}$ and given market maker belief $\widehat{D}$, the first-order condition for the insider’s problem (27) in the canonical game is
[…]
where $e_{k}$ denotes the $k$-th standard basis vector, and the expectation $\mathbb{E}^{(\hat{d};\widehat{D})}$$\cdot\,$] is taken over the distribution of $\hat{\omega}$ induced by the insider’s choice $\hat{d}$ when the market maker holds belief $\widehat{D}$. The terms in the first-order condition (30) are the isomorphic counterparts to those in Theorem 5.1.
6.2 Symmetric Equilibrium
The canonical game is symmetric—up to permutation on $k$, the game is identical conditional on each signal $s_{k}$. This symmetry suggests an equilibrium structure of the form
$D^{*}=\alpha{\bf Q},\mbox{ for some }\alpha>0$ (31)
where
[…]
This means that, conditional on $s_{k}$, the insider buys $\alpha(1-\frac{1}{K})$ shares of pseudo-security $k$ and sells the other pseudo-securities for $\frac{\alpha}{K}$ shares each. This demand corresponds to the $k$-th column of $D^{*}$ in (31). By substituting this equilibrium postulate into the first-order condition (30) conditional on each signal $s_{k}$ and collating the resulting vector equations side-by-side, we derive a matrix equation
$\Phi(\alpha){\bf Q}=0$ (33)
for some scalar function $\Phi\colon[0,\infty)\rightarrow\mathbb{R}$. We prove that each first-order condition is sufficient for optimality at such a fixed point of the collated first-order conditions (33). Thus, an equilibrium corresponds to a solution $\alpha^{*}>0$ to the equation $\Phi(\alpha)=0$. This is Theorem 6.5.

### v39

@@ 4.2 Sufficient Statistic
Given the market maker’s belief $\widetilde{W}(\,\cdot\,|s_{k})$ about insider demand conditional on $s_{k}$:
(i) The precision-weighted alignment score of order flow $\omega$ with $\widetilde{W}(\,\cdot\,|s_{k})$ is
[…]
(ii) The precision-weighted impact exposure of an insider portfolio $W(\,\cdot\,)$ to $\widetilde{W}(\,\cdot\,|s_{k})$ is
@@
5 Cross-Market Price Impact
Just as the market maker’s inference is infinite-dimensional, so is the insider’s portfolio choice. After observing his signal, the insider chooses an unrestricted portfolio $W(\,\cdot\,)$ of traded claims. Because there is no restriction on $W$, there is likewise no restriction on the pattern of order flow the market maker might rationalize under his belief. Prices are then set by how the realized order flow aligns with these conjectured patterns (see Theorem 4.1). As a result, changing $W$ changes the alignment of aggregate order flow with each signal, and therefore can move prices across all claims. Characterizing this price-impact requires taking marginal perturbations of the infinite-dimensional choice variable $W(\,\cdot\,)$.
Given market maker belief $\widetilde{W}$ and signal $s_{k}$, the insider’s portfolio choice problem (4) can be written as
@@ 5.1 First-Order Condition
[…]
@@
For any portfolio $W$ and any marginal trade package $v$, the insider’s marginal gain decomposes as
[…]
where
@@
Therefore, an optimal portfolio $W$ must satisfy the first-order condition
$\partial\!\kern 0.80002ptJ_{p}(v;W)\;=\;\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!AD}(v;W)\;+\;\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)\quad\mbox{for all marginal trade packages $v$.}$ (19)
The first-order condition (19) says that, at an optimal portfolio $W$, no marginal trade $v$ yields a positive gain: its payoff is exactly offset by its execution cost. This execution cost comprises, first, the expected execution-price term (17) and, second, the cross-market price-impact term (18).
  • Marginal Payoff $\partial\!\kern 0.80002ptJ_{p}(v;W)$ (16). This is the expected payoff of the incremental trade package $v$ under the signal-implied state probabilities $\eta(\,\cdot\,\mid s_{k})$. For example, if $v$ is an options package across strikes, then $\partial\!\kern 0.80002ptJ_{p}$ is the valuation of that package under the trader’s private distributional view (volatility, skew, tail risk, etc.).
@@
Single-Contract Cross Impact
The general price-impact term $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)$ is written for an arbitrary trade package $v$. The primitive empirical object is the cross-impact kernel—how signed order flow in one claim moves the price of another. Corollary 5.2 below isolates this kernel by specializing to the case where $v$ consists of one share of claim $x_{j}$. In that case, its impact exposure collapses to $\Pi_{insider}(v,\,\cdot\,;\widetilde{W})=\widetilde{W}(x_{j}\mid\,\cdot\,)/\sigma_{j}^{2}$—that is, how much a marginal trade in $x_{j}$ tilts the market maker’s posterior toward a given signal is exactly how much $x_{j}$ the market maker expects the insider to demand under that signal, adjusted for noise.
@@
Under the market maker’s belief $\widetilde{W}$, the price impact of marginal order flow for $x_{j}$ on the price of another claim $x_{i}$ given realized aggregate order flow $\omega$—is
@@
Averaging (20) over the distribution of realized order flow generated by portfolio $W$ gives the expected cross impact
@@
The intuition behind Corollary 5.2 is the same information-leakage slippage embedded in the general price-impact term $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)$, now in its sharpest single-contract form. A marginal trade in $x_{j}$ moves the price of $x_{i}$ to the extent that the information stories the market maker reads into $x_{j}$ order flow correlate with the stories that drive $x_{i}$’s payoff. This is exactly what the posterior covariance in (20) captures:
${\operatorname{\mathbb{Cov}}}\!\left(\eta(x_{i}|\,\cdot\,),\,\widetilde{W}(x_{j}|\,\cdot\,)\,\bigg|\,\omega\right).$
One can read $\widetilde{W}(x_{j}\mid\,\cdot\,)$ as the market maker’s story-by-story “informed-demand template” for $x_{j}$ and $\eta(x_{i}\mid\,\cdot\,)$ as $x_{i}$’s payoff profile across the same stories.
@@ 5.2 No-Arbitrage
[…]
$\Pi_{insider}(W,\,\cdot\,;\widetilde{W})=0,$
i.e., $W$ is orthogonal to the market maker’s belief about informed demand. These are “stealth” portfolios—they sit in the null space of the cross-impact kernel, so scaling them does not tilt the market maker’s posterior and therefore does not move prices. If such a zero-impact portfolio yielded nonzero payoff to the insider, the insider could lever it up without moving prices, generating an arbitrarily large payoff. This is Theorem 5.3.
@@
(No free lunch with zero price impact) Under market maker belief $\widetilde{W}$, an optimal insider portfolio can exist only if every zero price impact portfolio yields zero expected payoff to the insider. Otherwise, the insider can scale a zero-impact trading direction without moving prices and obtain arbitrarily large expected payoff.
@@ 6 Equilibrium
Up to this point we have solved each side of the market taking the other as given: the market maker’s pricing kernel as a mapping from cross-market order flow to the cross section of claim prices (Section 4), and the insider’s optimal portfolio in the same rich claim space (Section 5). In equilibrium, the conjectured “informed-demand template” that drives pricing must coincide with the insider’s optimal trades under that very pricing kernel—a joint fixed point of two high-dimensional problems. A key insight is that the trading game is isomorphic to a canonical game that is invariant to the information structure and noise intensities. [^15] This canonical-game reduction is fully robust. It holds whether the signal and state spaces are discrete or continuous. We keep the exposition in the finite-$S$ case here. Fully rigorous arguments for the continuous case are provided in Keller and Tseng (2026). This sets up a striking reduction—equilibrium, an infinite-dimensional fixed point a priori, is characterized by a single endogenous constant.
@@
(w.l.o.g.)   The market maker has a uniform prior on the signal space $S$.
For expositional convenience, we maintain Assumption 1 in this section. It is without loss of generality and serves only to streamline the derivations. Starting from any prior $\pi_{0}$, we can rescale payoffs and work as if the prior were uniform on $S$ by replacing $\eta(\,\cdot\,|s_{k})$ with $\frac{\pi_{0}(s_{k})}{1/|S|}\,\eta(\,\cdot\,|s_{k})$ for each $s_{k}\in S$. Once we characterize equilibrium under the uniform prior, comparative statics for arbitrary priors follow immediately. For instance, if two securities both yield higher payoffs under some $s_{k}$ compared to the other signals, then their cross price impact increases with the prior probability of $s_{k}$. For instance, if two securities tend to pay off together under some $s_{k}$, then putting more prior weight on $s_{k}$ makes their co-movement in order flow more diagnostic, so their equilibrium cross price impact increases with $\pi_{0}(s_{k})$. This generalizes from the toy model of Section 2; see footnote 9.
@@ 6.1 Canonical Game
In equilibrium, the portfolios that yield zero expected payoff to the insider must be exactly the portfolios that generate zero price impact. Zero payoff implies zero price impact because equilibrium beliefs are correct (Definition 1). Conversely, by Theorem 5.3, a zero-impact portfolio with nonzero payoff would be a stealth arbitrage that the insider could scale without moving prices.
This coincidence translates to a restriction on the relevant trading subspace. The insider’s expected payoff is driven only by how $W$ loads on the signal-implied payoff distributions $\{\eta(\,\cdot\,\mid s_{k})\}_{k}$. Any portfolio $W$ that is orthogonal to each distribution—i.e., $\sum_{i}W(x_{i})\eta(x_{i}\mid s_{k})=0$ for all kk—has zero expected payoff under every signal. Hence, without loss, equilibrium informed demand can be taken to lie in the linear span of $\{\eta(\,\cdot\,\mid s_{k})\}_{k}$. Conversely, to rule out stealth arbitrage, the market maker’s (noise-adjusted) equilibrium “informed-demand templates” must cover every payoff-relevant direction—otherwise the insider could hide profits in a trading direction the market maker cannot see. That is, each $\eta(\,\cdot\,\mid s_{k})$ must lie in the span of $\{W^{*}(\,\cdot\,\mid s_{\ell})/\sigma_{\,\cdot\,}^{2}\}_{\ell}$. We record these two spanning restrictions in Proposition 6.1.
@@
(ii) Conversely, to prevent arbitrage, each $\eta(\,\cdot\,|s_{k})$ must lie in the linear span of $\{\frac{W^{*}(\,\cdot\,|s_{l})}{\sigma_{\,\cdot\,}^{2}}\}_{l}$.
To make the reduction in Proposition 6.1 explicit, we parameterize portfolios and beliefs on the payoff-relevant span of $\{\eta(\,\cdot\,\mid s_{k})\}_{k=1}^{K}$ by their loadings on these signal-implied payoffs. Accordingly, write any portfolio in this span as
@@
Likewise, expand the market maker’s signal-contingent “informed-demand template” $\widetilde{W}(\,\cdot\,\mid s_{k})$ in the same basis,
$\widetilde{W}(\,\cdot\,\mid s_{k})=\sum_{l=1}^{K}\tilde{d}^{(k)}_{l}\,\eta(\,\cdot\,\mid s_{l}),\qquad 1\leq k\leq K,$
and stack the loading vectors as columns to form the belief matrix
$\widetilde{D}\;=\;[\tilde{d}^{(1)}\ \cdots\ \tilde{d}^{(K)}]\in\mathbb{R}^{K\times K}.$
@@
The information intensity matrix ${\bf L}\in\mathbb{R}^{K\times K}$ of the trading game is the positive-semidefinite matrix satisfying
${\bf L}^{2}=\left[\sum_{i}\frac{\eta(x_{i}\mid s_{k})\,\eta(x_{i}\mid s_{l})}{\sigma_{i}^{2}}\right]_{1\leq k,l\leq K},\qquad{\bf L}^{T}={\bf L}.$
The information intensity matrix $\bf L$} is the cross-market analogue of the signal-to-noise ratio $\sigma_{v}/\sigma_{\varepsilon}$ from the single-asset Kyle setting (6). It aggregates the (noise-adjusted) co-variations of the signal-implied payoff profiles $\{\eta(\,\cdot\,\mid s_{k})\}_{k}$ across claims, so a larger $\bf L$} means that cross-market order flow is more diagnostic of which information story is active. [^16] Here, “larger” is in the positive-semidefinite order. This order is partial—unlike the single-asset setting, information intensity is not always directly comparable between two specifications.
Conversion to Information Units
Accordingly, $\bf L$} provides exactly the right rescaling to convert the raw loadings in $d$ and $\widetilde{D}$ into information units: premultiplying by $\bf L$} “whitens” the payoff-relevant span so that one unit of loading corresponds to one unit of noise-adjusted informational exposure. We therefore pass to the information-normalized loading vector $\hat{d}$ and belief matrix $\widehat{D}$ given by
$\hat{d}$ $=\;{\bf L}d,$ (22)
$\widehat{D}$ $=\;{\bf L}\widetilde{D}\;=\;[\hat{d}^{(1)}\ \cdots\ \hat{d}^{(K)}].$
Canonical Game
The information-unit conversion (22) strips away the idiosyncrasies of the payoff distributions and noise intensities and isolates the strategic core of the trading problem. In information units, the payoff-relevant span becomes orthonormal—each trading direction carries one unit of signal relative to noise. The trading problem can then be read as trading orthogonal information contracts—one per information story—under unit noise intensity. The next theorem records this equivalence.
@@
(Canonical Game - Informal Statement) The original trading game is isomorphic to the canonical game in which:
  • the traded claims are information contracts—one for each signal $s_{k}$, paying $1$ under story $k$ and 0 otherwise;
  • noise trading is i.i.d. $\mathcal{N}(0,1)$ across contracts;
  • competitive prices are the market maker’s posterior probabilities over the signals.
The full formal statement is deferred to Appendix Section A.5.
Equilibrium in the Canonical Game
An equilibrium in the canonical game is summarized by a $K\times K$ matrix $D^{*}=[\delta^{(1)}\ \cdots\ \delta^{(K)}]$ where the $k$-th column $\delta^{(k)}\in\mathbb{R}^{K}$ specifies the insider’s demand for the information contracts when story $k$ is active. Given the market maker’s belief $D^{*}$, prices in the canonical game are posterior probabilities over the $K$ stories, and each $\delta^{(k)}$ is optimal under those very prices. In other words, $D^{*}$ is the fixed point that closes the canonical game.
Mapping Back to the Original Game
Given a canonical-game equilibrium $D^{*}$, we reverse the information-unit conversion (22) to express the same equilibrium trades as loadings on the original payoff distributions $\{\eta(\,\cdot\,\mid s_{l})\}_{l=1}^{K}$. Define the canonical form loadings matrix [^17] If $\bf L$} is singular, interpret ${\bf L}^{-1}$ as the Moore–Penrose pseudoinverse.
${\bf L}^{-1}D^{*}\;=\;[\beta^{(1)}\ \cdots\ \beta^{(K)}]\in\mathbb{R}^{K\times K},$ (23)
so the $k$th column $\beta^{(k)}$ collects the insider’s equilibrium loadings across payoff distributions when signal $s_{k}$ is realized. The corresponding informed demand is then
$W^{*}(\,\cdot\,\mid s_{k})\;=\;\sum_{l=1}^{K}\beta_{l}^{(k)}\,\eta(\,\cdot\,\mid s_{l}).$ (24)
6.2 Symmetry and One-Parameter Equilibrium
The canonical-game reduction sets up the central observation. Despite the trading game being infinite-dimensional, its equilibrium restriction is extremely tight—it collapses to a single endogenous scalar.
The canonical game treats the information stories symmetrically: relabeling stories just relabels markets. This symmetry forces a simple long-short equilibrium structure. Conditional on story $k$, the insider goes long the $k$-th information contract and finances it by shorting the average of the remaining contracts.
Let $\bf Q$} denote the $K\times K$ matrix that encodes this “buy the active story and sell the alternatives” strategy, column by column:
${\bf Q}\;=\;{\bf I}-\frac{1}{K}\,\bar{e}\,\bar{e}^{T},\qquad\mbox{ where }\bar{e}\;=\;(1,\ldots,1)^{T}\in\mathbb{R}^{K}.$ (25)
So the $k$-th column of $\bf Q$} puts portfolio weights $1-\tfrac{1}{K}$ on contract $k$ and $\tfrac{1}{K}$ on each of the remaining contracts. Accordingly, we parameterize candidate equilibria by a single parameter $\alpha>0$, with corresponding strategy matrices of the form
$D^{*}\;=\;\alpha\,{\bf Q}.$ (26)
Imposing insider optimality within this symmetric class reduces the fixed point condition to a single scalar equation in $\alpha$. Theorem 6.3 shows that it admits a solution $\alpha^{*}>0$, delivering equilibrium strategy matrix $D^{*}=\alpha^{*}{\bf Q}$.